OpenStudy (anonymous):

write an equation for the linear function f satisfying the given conditions. f(-5)=-1 and f(2)=4

OpenStudy (anonymous):

y=mx+b -1=m(-5)+b 4=m(2) + b set b = b -5m-1 = -2m + 4 -3m = 5 m=-5/3

OpenStudy (anonymous):

you would rewrite it in ax+by=c or if they want slope intercept form you would write it in y=mx+b

OpenStudy (anonymous):

b=2/3 so y=-5/3x + 2/3

OpenStudy (anonymous):

do you understand sdejayo?

OpenStudy (anonymous):

Yeah. Thanks!